OpenStudy (anonymous):

can someone show/explain how to do the following problem: 4 (1/2+1/4)

OpenStudy (anonymous):

\[4\left(\frac{1}{2}+\frac{1}{4}\right)=4\times \frac{1}{2}+4\times \frac{1}{4}=2+1\]

OpenStudy (anonymous):

thank you so much @satellite73 :)